Home‎ > ‎Crew 162‎ > ‎

0126-DailySummaryReport

DRS Daily Summary Report for Sol 03, Crew 162

Summary Title: Do not go gentle into that good night. Rage, rage against
the dying of the light.

Mission Status:  Our third Sol on Mars was a model day. It had structure
and purpose and the team is functioning as a well-oiled machine. We have
a schedule that works and a philosophy of helping whenever and wherever
needed to support one another. Research filled our afternoon and we are
excited for our upcoming media visit on Thursday (Sol 05).

Sol Activity Summary: Team rose early for a stargazing opportunity prior
to sunrise but was not able to catch the Iridium satellite pass-by or
get the telescope fired up in time. Still, we had a good time bundling
up against the cold morning air and trying our best to capture digital
photos of the planets, moon, hab and sunrise from the observatory
tunnel. During our morning planning conference, we defined our
short-term team goals and got ourselves properly organized for the week.
We have much work ahead of us but are ready to accept the challenges of
life in sim. We trained on EVA prep and HabCom skills as well as how to
perform engineering checks. Research activities were the focus of the
afternoon while we watched the movie Interstellar, a favorite of our
team who ponder things like planetary stability and the search for
Earth-like planets in habitable zones around other stars that we might
one day colonize. Our EVA for hunting micrometeorites went off without
any issues and we will see in a few days if the first batch of topsoil
collected by the Commander and HSO revealed any particles from outer
space.

Look Ahead Plan: We are waiting on a response to our questions about the
media visits this week and next. We will continue our research projects
and report on our progress. We are excited about the upcoming
opportunities to discuss not only our research and graduate school
research but to help showcase the best features of the MDRS program at
this site. We will be sprucing up the hab in anticipation of our
visiting crew members' docking.

Anomalies in work:  No problems experienced on Sol 03

Weather: Clear, cool and beautiful day today on Mars.

Crew Physical Status: Crew is in good health but feeling the effects of
a very full day today.

EVA: Commander and HSO completed a short (30 minute) on-foot EVA near
the hab to collect topsoil samples that may contain micrometeorites that
have fallen through the atmosphere. No issues encountered. Team will
submit an EVA summary during tonight's comm window.

No EVA requested for tomorrow, Sol 04. We would like to perform an EVA
during the media visit on Thursday, Sol 05.

Reports to be filed tonight:
-Daily Summary
- Commander’s Report
-Green Hab Report
-EVA Summary
-Engineering Report
-Journalist Report
- Daily Photos

Support Requested: No support requested beyond the engineering resource
requests of water and propane refills and trash removal.